OpenStudy (anonymous):

how do i find Find two arithmetic means between 67 and 90.

OpenStudy (asnaseer):

call the two unknowns a and b. you then have a sequence: 67, a, b, 90

OpenStudy (asnaseer):

the difference between pair of terms has to be the same

OpenStudy (amistre64):

67 m1 m2 90 asaneer got it :)

OpenStudy (asnaseer):

use that to form a set of equations and solve to find a and b

OpenStudy (amistre64):

midpoint is 1 mean between 2 values: (a+b)/2 = common difference in order to find 2 means, we divide by 3: (a+b)/3 = common difference in order to find n means, we divide by n+1: (a+b)/(n+1) = common difference

OpenStudy (amistre64):

...or something akin to that eh :)

OpenStudy (asnaseer):

@maine2 do you know how to proceed?

OpenStudy (asnaseer):

another way of viewing this is to take a standard arithmetic sequence: a, a+d, a+2d, a+3d, ... where a is the first term and d is the common difference. in your case you have 67, m1, m2, 90 so we know a=67 and a+3d=90 which implies 67 + 3d = 90 use that to find d and then you have the full sequence

OpenStudy (anonymous):

how do find d

OpenStudy (asnaseer):

solve this: 67 + 3d = 90

OpenStudy (anonymous):

23

OpenStudy (asnaseer):

please show your steps so that I can spot where you may have made a mistake

OpenStudy (anonymous):

i just dif 90-67=23

OpenStudy (asnaseer):

start with: 67 + 3d = 90 and show each step please

OpenStudy (asnaseer):

?

OpenStudy (anonymous):

67+3d=90 -67 -67 3d=23 divide by 3 d=7.66

OpenStudy (asnaseer):

correct, although 7.66 is an approximation. I would leave it as a fraction, i.e. \(\displaystyle7\frac{2}{3}\)

OpenStudy (asnaseer):

now recall that the sequence you had was of the form: a, a+d, a+2d, a+3d and you know a=67 and you just calculated d=7 2/3, so now you should be able to find the middle two terms

OpenStudy (anonymous):

ok so a+d is 234/3

OpenStudy (asnaseer):

don't keep it as a mixed fraction, instead work it out as a+d=67 + 7 2/3 = ?

OpenStudy (asnaseer):

67 + 7 = ?

OpenStudy (anonymous):

74

OpenStudy (asnaseer):

sorry I meant to say "don't keep it as an improper fraction.." up there.

OpenStudy (asnaseer):

yes, so a+d = 67 + 7 2/3 = 74 2/3

OpenStudy (asnaseer):

now you need a+2d = (a+d)+d = 74 2/3 + 7 2/3 = ?

OpenStudy (anonymous):

so it is 81

OpenStudy (asnaseer):

74 + 7 = 81 yes but you also need the fractional part: 2/3 + 2/3 = ?

OpenStudy (anonymous):

1.3

OpenStudy (asnaseer):

:/ yes - but leave as a fraction please, so its 1 1/3

OpenStudy (asnaseer):

so 3rd term = 81 + 1 1/3 = ?

OpenStudy (anonymous):

so it is 82 1/3

OpenStudy (asnaseer):

yes, so your final sequence is:\[67, 74\frac{2}{3},82\frac{1}{3},90\]

OpenStudy (asnaseer):

well done! :)

OpenStudy (anonymous):

thanks asnaseer

OpenStudy (asnaseer):

yw :)